OpenStudy (anonymous):

A certain oil well that yields 400 barrels per month will run dry in 2 years. Supposing that the current price of crude oil is 98 per barrel and that the price of crude oil will rise at a constant rate of 40 cents per month over the next two years, how much revenue will be obtained from the well over the next two year

OpenStudy (anonymous):

$946560

OpenStudy (anonymous):

Let x = # of months in the future from 0 to 24. revenue = 24 ∫ 400(95 + 0.3x) dx = 0 24 ∫ (38000 + 120x) dx = 0 .......................24 38000x + 60x^2 | = .......................0 38000(24) + 60(24^2) - (38000(0) + 60(0^2)) = $946560